\begin{thebibliography}{10}

\bibitem{ADNEY_YEN:1965}
J.~E. Adney and T.~Yen, \emph{Automorphisms of a {$p$}-group}, Illinois Journal
  of Mathematics \textbf{9}, 137--143, 1965.

\bibitem{BARR:1971}
M.~Barr, \emph{Exact categories}, in: Lecture Notes in Mathematics
  \textbf{236}, 1--120, 1971.

\bibitem{BORCEUX_CLEMENTINO:2005}
F.~Borceux and M.~M. Clementino, \emph{Topological semi-abelian algebras},
  Advances in Mathematics \textbf{190}(2), 425--453, 2005.

\bibitem{BORCEUX_BOURN:2007}
F.~Borceux and D.~Bourn, \emph{Split extension classifier and centrality},
  Contemporary Mathematics, Categories in algebra, geometry and mathematical
  physics \textbf{431~}, American Mathematical Society, Providence, RI,
  85--104, 2007.

\bibitem{BORCEUX_JANELIDZE_KELLY:2005a}
F.~Borceux, G.~Janelidze, and G.~M. Kelly, \emph{Internal object actions},
  Commentationes Mathematicae Universitatis Carolinae \textbf{46}(2), 235--255,
  2005.

\bibitem{BORCEUX_JANELIDZE_KELLY:2005b}
F.~Borceux, G.~Janelidze, and G.~M. Kelly, \emph{On the representability of
  actions in a semi-abelian category}, Theory and Applications of Categories
  \textbf{14}(11), 244--286, 2005.

\bibitem{BOURN:1991}
D.~Bourn, \emph{Normalization equivalence, kernel equivalence and affine
  categories}, Lecture Notes in Mathematics, Category theory ({C}omo, 1990)
  \textbf{1488~}, Springer, Berlin, 43--62, 1991.

\bibitem{BOURN:2013b}
D.~Bourn, \emph{Two ways to centralizers of equivalence relations}, Applied
  categorical structures \textbf{22}(5-6), 833--856, 2014.

\bibitem{BOURN_GRAN:2002}
D.~Bourn and M.~Gran, \emph{Central extensions in semi-abelian categories}
  \textbf{175~}, 31--44, 2002, Special volume celebrating the 70th birthday of
  Professor Max Kelly.

\bibitem{BOURN_JANELIDZE:1998}
D.~Bourn and G.~Janelidze, \emph{Protomodularity, descent and semidirect
  products}, Theory and Applications of Categories \textbf{4}(2), 37--46, 1998.

\bibitem{BOURN_JANELIDZE:2009}
D.~Bourn and G.~Janelidze, \emph{Centralizers in action accessible categories},
  Cahiers de Topologie et G\'eom\'etrie Diff\'erentielles Cat\'egoriques
  \textbf{50}(3), 211--232, 2009.

\bibitem{CAGLIARI_CLEMENTINO:2019}
F.~Cagliari and M.~M. Clementino, \emph{Topological groups have representable
  actions}, Bulletin of the Belgian Mathematical Society. Simon Stevin
  \textbf{26}(4), 519--526, 2019.

\bibitem{CIGOLI_MANTOVANI:2012}
A.~Cigoli and S.~Mantovani, \emph{Action accessibility via centralizers},
  Journal of Pure and Applied Algebra \textbf{216}(8-9), 1852--1865, 2012.

\bibitem{CIGOLI_GRAY_VAN_DER_LINDEN:2015a}
A.~S. Cigoli, J.~R.~A. Gray, and T.~Van~der Linden, \emph{On the normality of
  {H}iggins commutators}, Journal of Pure and Applied Algebra \textbf{219}(4),
  897--912, 2015.

\bibitem{CIGOLI_MONTOLI:2015}
A.~S. Cigoli and A.~Montoli, \emph{Characteristic subobjects in semi-abelian
  categories}, Theory and Applications of Categories \textbf{30}, 206--228,
  2015.

\bibitem{CIGOLI_GRAY_VAN_DER_LINDEN:2015b}
A.~S. Cigoli, J.~R.~A. Gray, and T.~{Van der Linden}, \emph{Algebraically
  coherent categories}, Theory and Applications of Categories \textbf{30}(54),
  1864--1905, 2015.

\bibitem{GRAN_GRAY:2021}
M.~Gran and J.~R.~A. Gray, \emph{Action representability of the category of
  internal groupoids}, Theory and Applications of Categories \textbf{37}, Paper
  No. 1, 1--13, 2021.

\bibitem{GRAN_STERCK_VERCRUYSSE:2019}
M.~Gran, F.~Sterck, and J.~Vercruysse, \emph{A semi-abelian extension of a
  theorem by takeuchi}, Journal of Pure and Applied Algebra, 2019.

\bibitem{GRAY:2013b}
J.~R.~A. Gray, \emph{Normalizers, centralizers and action representability in
  semi-abelian categories}, Applied Categorical Structures \textbf{22}(5-6),
  981--1007, 2014.

\bibitem{GRAY:2022}
J.~R.~A. Gray, \emph{Complete objects in categories}, Journal of Pure and
  Applied Algebra \textbf{226}(4), Paper No. 106857, 26, 2022.

\bibitem{GRAY_VAN_DER_LINDEN:2015}
J.~R.~A. Gray and T.~V. der Linden, \emph{Peri-abelian categories and the
  universal central extension condition}, Journal of Pure and Applied Algebra
  \textbf{219}(7), 2506--2520, 2015.

\bibitem{HUQ:1968}
S.~A. Huq, \emph{Commutator, nilpotency and solvability in categories},
  Quarterly Journal of Mathematics \textbf{19}(1), 363--389, 1968.

\bibitem{JANELIDZE_MARKI_THOLEN:2002}
G.~Janelidze, L.~M\'arki, and W.~Tholen, \emph{Semi-abelian categories},
  Journal of Pure and Applied Algebra \textbf{168}, 367--386, 2002.

\bibitem{ORZECH:1972}
G.~Orzech, \emph{Obstruction theory in algebraic categories \uppercase{I, II}},
  Journal of Pure and Applied Algebra \textbf{2}(4), 287--314 and 315--340,
  1972.

\end{thebibliography}

\documentclass{article}
\usepackage{amsmath,amsthm,amsfonts,amssymb,xcolor,enumerate,stmaryrd,mathtools}

\usepackage[all]{xy}

\newcommand{\C}{\mathbb{C}}

\newcommand{\Aut}{\textnormal{Aut}}
\newcommand{\Der}{\textnormal{Der}}

\newcommand{\SplExt}{\textbf{SplExt}}
\newcommand{\bwnh}{\textbf{(WNH)}}
\newcommand{\bnh}{\textbf{(NH)}}
\newcommand{\wnh}{\textnormal{(WNH)}}
\newcommand{\nh}{\textnormal{(NH)}}
\newcommand{\comm}[2]{\gamma_{[#1, #2]}}
\newcommand{\como}[2]{[#1, #2]}
\newcommand{\spexc}[1]{\llbracket#1\rrbracket}

\newcommand{\ke}{\textnormal{ker}}
\newcommand{\coke}{\textnormal{coker}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{example}[theorem]{Example}
\newtheorem{remark}[theorem]{Remark}

\begin{document}
\title{A note on split extension classifiers of perfect objects}
\author{J. R. A. Gray}
\maketitle
\begin{abstract}
We show that for a pointed protomodular category $\C$ 
satisfying a certain condition on those Huq commutators which exist, if $X$ is a
perfect object in $\C$ such
that
the split extension classifier $\spexc{X}$ exists, then the centralizer of
the \emph{conjugation} morphism
$c_X : X\to \spexc{X}$ is trivial and hence $\spexc{X}$ has trivial center.
\end{abstract}
\section{Introduction and preliminaries}
Recall that the automorphisms of a group commuting with every inner automorphism
are called central. For a group $X$ these automorphisms form a normal subgroup
$\Aut_c(X)$ of the automorphism group $\Aut(X)$ which has been the subject of
several papers whose study goes back at least to J.~E.~Adney and T.~Yen
\cite{ADNEY_YEN:1965}.
One of the methods of studying these automorphisms is via the so called
\emph{Adney-Yen}
map which is an injective map from $\Aut_c(X)$ to the set
of homomorphisms from $X$ to its center $Z(X)$.
As a immediate consequence of the existence of this injective map
one sees
that if $X$ is perfect, then $\Aut_c(X)$ is trivial.

Our main aim in
this paper is to generalize this result to, in particular,
arbitrary \emph{semi-abelian action representable} categories satisfying
a mild condition on commutators. The terminology will be recalled below.

As a special case we will
recover the following known fact about Lie algebras.
For a fixed commutative ring $R$
and a Lie algebra $X$ 
over $R$, the Lie algebra of derivations of $X$, $\Der(X)$, has underlying abelian group
the subgroup of $R$-linear maps $d:X\to X$ satisfying $d([x,y])=[d(x),y]+[x,d(y)]$
for each $x,y$ in $X$, and has Lie bracket defined by $[d,e]=d\circ e-e\circ d$.
For each $x$ in $X$, the map $y\mapsto [x,y]$ determines a derivation called
an inner derivation.
In analogy with the group theoretic case (in fact both
are special cases of the same categorical construction) derivations which
commute with (annihilate)
every inner derivation are called central and form an ideal of $\Der(X)$ which
we denote by $\Der_c(X)$. The analogous fact is that
if $X$ is perfect, then $\Der_c(X)$ is trivial.

We now recall the necessary background and introduce notation
so that we can precisely state our main aim.

For a pointed category $\C$ we will denote by $0$ a chosen
zero object (that is, an object which is both initial and terminal). For
objects $A$ and $B$ in $\C$ we will also denote by $0$
the unique morphism from $A$ to $B$ which factors through the zero
object $0$. The product of a pair of objects $A$ and $B$ will be written
$A\times B$ with first and second projection denoted by $\pi_1$ and $\pi_2$,
respectively. The coproduct will be written $A+B$ with coproduct inclusions
$\iota_1$ and $\iota_2$. For morphisms $u : W \to A$, $v: W\to B$, $f:A\to C$
and $g : B\to C$
we denote by $\langle u,v\rangle : W\to A\times B$ and
$[f,g]:A+B\to C$ the unique morphisms with
$\pi \langle u, v\rangle=u$, $\pi_2 \langle u,v\rangle = v$, $[f,g]\iota_1=f$
and $[f,g]\iota_2 = g$.

Recall that a split extension in a pointed category $\C$ is a
diagram
\begin{equation}
\label{diagram:split}
\vcenter{
\xymatrix{
X \ar[r]^{\kappa} & A\ar@<0.5ex>[r]^{\alpha} & B\ar@<0.5ex>[l]^{\beta}
}
}
\end{equation}
in $\C$ where $\alpha \beta = 1_B$ and $\kappa$ is the kernel of
$\alpha$. A morphism of split extensions in $\C$ is a diagram
\begin{equation}
\label{diagram:morphism_of_split}
\vcenter{
\xymatrix{
X \ar[r]^{\kappa}\ar[d]^{f} & A\ar@<0.5ex>[r]^{\alpha}\ar[d]^{g} &
B\ar@<0.5ex>[l]^{\beta}\ar[d]^{h}\\
Z\ar[r]_{\sigma} & C\ar@<0.5ex>[r]^{\gamma} & D\ar@<0.5ex>[l]^{\delta}
}
}
\end{equation}
such that the top and bottom rows are split extensions (the domain
and codomain, respectively) and $g\kappa = \sigma f$, $g\beta = \delta h$
and $h\alpha = \gamma g$. We denote by $\SplExt(\C)$ the category of
split extensions in $\C$, and by $K$ and $P$ the functors sending \eqref{diagram:split}
to $X$ and $B$ respectively,
and \eqref{diagram:morphism_of_split} to $f$ and $h$ respectively.
We denote by $K^{-1}(X)$ the fiber of the functor $K$ above $X$, that is
the subcategory of $\SplExt(\C)$ with objects and morphism mapped by $K$
to $X$ and $1_X$, respectively. Following F.\ Borceux, G.\ Janelidze and
G.\ M.\ Kelly
\cite{BORCEUX_JANELIDZE_KELLY:2005a}
 we call a terminal object in $K^{-1}(X)$ a generic split
extension with kernel $X$ which we will denote by
\begin{equation}
\label{diagram:generic_split}
\vcenter{
\xymatrix{
 X \ar[r]^-{k} & \spexc{X}\ltimes X \ar@<0.5ex>[r]^-{p_1} & \spexc{X}.\ar@<0.5ex>[l]^-{i}
}
}
\end{equation}
If $X$ is an object in a pointed category $\C$ admitting a generic
split extension with kernel $X$, then we will denote by $c_X : X\to \spexc{X}$
the morphism forming part of the unique morphism
\begin{equation}
\label{diagram:conj}
\vcenter{
\xymatrix{
X \ar[r]^-{\langle 0,1\rangle} \ar@{=}[d]
& X\times X \ar@<0.5ex>[r]^-{\pi_1}\ar[d]^{\phi}
&  X \ar@<0.5ex>[l]^-{\langle 1,1\rangle}\ar[d]^{c_X}\\
X \ar[r]^-{k} & \spexc{X}\ltimes X \ar@<0.5ex>[r]^-{p_1} & \spexc{X}\ar@<0.5ex>[l]^-{i}
}
}
\end{equation}
in $K^{-1}(X)$. Recall that there is also morphism
$p_2 : \spexc{X}\ltimes X\to \spexc{X}$ which together with $p_1$ and $i$ form part of
a canonical groupoid structure (see
\cite{BORCEUX_BOURN:2007}
or the preliminaries of
\cite{GRAY:2022}),
and furthermore $c_X = p_2 k$.
Note that if $\C$ is the category of groups, then $\spexc{X}=\Aut(X)$
and $c_X$ is the homomorphism sending each $x$ in $X$ to the inner
automorphism determined by conjugation 
(that is $x \mapsto (y \mapsto xyx^{-1})$) \cite{BORCEUX_JANELIDZE_KELLY:2005b}.
On the other hand if $\C$ is the category of Lie algebras over a
commutative ring, then $\spexc{X}=\Der(X)$ and $c_X$ is the homomorphism
sending each $x$ in $X$ to the inner derivation determined by $x$ (that is
$x\mapsto (y\mapsto [x,y])$) \cite{BORCEUX_JANELIDZE_KELLY:2005b}.

Recall that a pointed finitely complete category can equivalently be defined
to be (Bourn)-protomodular
\cite{BOURN:1991}
if it satisfies the split short five lemma, that is, for
each morphism of split extensions \eqref{diagram:morphism_of_split} if $f$ and
$h$ are isomorphisms, then $g$ is an isomorphism. A category is regular if
it has all finite limits, regular epimorphisms are pullback stable in it, and
the coequalizer of kernel pairs exist in it. A category is homological 
if it is a regular pointed protomodular category.
A pair of
morphisms $f:A \to C$ and $g: B\to C$ in a pointed protomodular category
$\C$ with binary products (Huq)-commute \cite{HUQ:1968} if there exists a morphism
$\varphi : A\times B \to C$ making the diagram
\[
\xymatrix{
A\ar[r]^-{\langle 1,0\rangle}\ar[dr]_{f} &
A\times B \ar[d]^{\varphi}
& B\ar[l]_-{\langle 0,1\rangle}\ar[dl]^{g}\\
& C &
}
\]
commute. For $f$ and $g$ as above we use the following (equivalent) definition
of the Huq-commutator (originally defined in \cite{HUQ:1968}).
The Huq commutator
$\comm{f}{g} : \como{(A,f)}{(B,g)} \to C$ of $f$ and $g$ is
initial amongst those normal monomorphisms $n$ with codomain $C$ such that
 $\coke(n)f$ and $\coke(n)g$ commute (where $\coke(n)$ is the cokernel of $n$). Note that we will write $\como{X}{X}$ instead of
$\como{(X,1_X)}{(X,1_X)}$.
The centralizer of a morphism $f: A\to C$ is the terminal object in
the category of morphisms which commute with $f$ and will be denoted
$z_f:Z_C(A,f)\to C$.
Recall also that an object
$X$ is commutative if $1_X$ commutes with $1_X$ and is perfect if the only
morphisms from it to commutative objects are zero morphisms.

Recall that a category $\C$ is (Barr)-exact
\cite{BARR:1971},
if it is regular, and equivalence relations
in $\C$ are effective (i.e.{} are the kernel pair of some morphism).
A pointed protomodular
category is semi-abelian
\cite{JANELIDZE_MARKI_THOLEN:2002}
if it has binary coproducts and is Barr-exact.
In agreement with the definitions in the semi-abelian context
(see \cite{GRAY_VAN_DER_LINDEN:2015} and \cite{CIGOLI_GRAY_VAN_DER_LINDEN:2015a},
respectively) we consider the following conditions on a pointed category $\C$:\\

\noindent\bwnh{}. For each normal monomorphism
$k : K\to X$ in $\C$ if the commutator $\comm{1_K}{1_K}$ exists,
then the morphism $k\comm{1_K}{1_K}$ is the commutator morphism of $k$ with itself.\\

\noindent\bnh{}. For monorphisms $m:X\to Y$ and $k:K\to X$ and $l:L\to X$ in $\C$
if $\comm{k}{l}$ exists and $mk$ and $ml$ are normal monomorphisms,
then $m\comm{k}{l}$ is the commutator morphism of $mk$ and $ml$.\\

It turns out that \nh{} and \wnh{} can be
rephrased in terms of Higgins commutators in the appropriate context where those
commutators exist (see \cite{GRAY_VAN_DER_LINDEN:2015} and \cite{CIGOLI_GRAY_VAN_DER_LINDEN:2015a} where, in fact,
\wnh{} and \nh{} were defined in terms of Higgins commutators).

Our main aim is to show that in a pointed protomodular category satisfying
\wnh{},
if $X$
is perfect and the generic split extension with kernel $X$ exists, then the
centralizer of $c_X$ is a zero morphism.

In addition we produce categorical explanations of the following facts:
\begin{enumerate}[(i)]
\item if $f:X\to Y$ is a central extension of groups, and $\cdot$ and
$*$ are actions of a group $B$ on $X$ such that, for each $b$ in $B$ 
each $x$ in $X$ we have that $f(b\cdot x)=f(b*x)$, then both actions
have the same restriction to $[X,X]$ (Proposition \ref{proposition:actions_2});
\item if $\theta:X\to X$ is a central automorphism, then for each
$x$ and $y$ in $X$ we have that $\theta([x,y])=[x,y]$ (Corollary \ref{corollary:a}).
\end{enumerate}

\section{Lifting extensions and properties of morphisms with kernel a central monomorphism}
In this section we establish some results about when split extensions can be
lifted and the uniqueness of such liftings, which we will need in final
section.

We begin by recalling some known facts.
It is well known and easy to prove for a central extension $f:X\to Y$ 
of groups and for $a, a', b$ and $b'$ in $X$
if  $f(a)=f(a')$, and $f(b)=f(b')$, then $[a,b]=[a',b']$. Categorically this
follows from the following fact
(which is essentially the same as Proposition 3.1 of \cite{BOURN_GRAN:2002}).
\begin{proposition}
\label{proposition:monomorphic behavior of central extensions}
Let $\C$ be a pointed protomodular category, $f:A\to C$ a morphism
in $\C$, and $k:K\to A$ the kernel of $f$. If $k$ is central, then
\begin{enumerate}[(a)]
 \item the pair $\varphi,\pi_2 : K\times A \to A$ (where $\varphi: K\times A\to A$ the
  cooperator of $k$ and $1_A$) is the
 kernel pair of $f$, and the composite $\langle 0,1\rangle \comm{1_A}{1_A}$ is the
  commutator morphism of $1_{K\times A}$ with itself (provided $\comm{1_A}{1_A}$ exists);
\item  For pair of morphisms $u,v : W\to A$ if $fu=fv$, then
 $u\comm{1_W}{1_W} = v\comm{1_W}{1_W}$ (provided $\comm{1_W}{1_W}$ exists).
\end{enumerate} 
\end{proposition}
We will also need the following slight strengthening of Lemma 2.6 of
\cite{CIGOLI_GRAY_VAN_DER_LINDEN:2015a}.
 \begin{lemma}
 \label{lemma:lifting}
  Let $\C$ be a pointed protomodular category.
  For each split extension as displayed in the bottom row of the diagram
  \[
   \xymatrix{
    X'\ar[d]_{u}\ar@{-->}[r]^{\kappa'}&
    A'\ar@{-->}@<2.15pt>[r]^{\alpha'}\ar@{-->}[d]^{v}&
    B\ar@{-->}@<2.15pt>[l]^{\beta'}\ar@{==}[d]\\
    X\ar[r]_{\kappa}&A\ar@<2.15pt>[r]^{\alpha}&B\ar@<2.15pt>[l]^{\beta}
   }
  \]
  there exists a unique (up to isomorphism) lifting along a monomorphism $u$ to a morphism of split
  extensions, as display with dotted arrows,
  as soon as $\kappa u$ is normal (or more generally Bourn-normal).
 \end{lemma}
\begin{proof}
The existence of a lifting follows from Lemma 2.6 of
\cite{CIGOLI_GRAY_VAN_DER_LINDEN:2015a}.
Uniqueness follows from e.g.~Lemma 4.1 of \cite{GRAY:2013b}.
\end{proof}

It is easy to check and well known that for a central extension $f:X\to Y$
of a group if $\cdot, *$ are actions of a group $B$ on $X$ 
such that for each $x\in X$ and for each $b\in B$ we have that
$f(b\cdot x) = f(b*x)$, then both actions have the same restriction to
$\como{X}{X}$. We will see that this can be recovered from
the following categorical fact (which is closely related to
Lemma 3.10 of \cite{GRAY_VAN_DER_LINDEN:2015}):
\begin{proposition}\label{proposition_similtanius_lifting}
Let $\C$ be a pointed protomodular category satisfying \wnh.
For a morphism $f: X\to Z$ with kernel a central monomorphism and
for morphisms of split extensions displayed in the diagram
 \begin{equation}
  \vcenter{
   \label{diagram:morphisms_of_split_exts_with_the_same_morphism_at_kernel}
  \xymatrix{
  X
  \ar[r]^{\kappa_1}
  \ar[d]_{f}
  &
  A_1
  \ar@<2.00pt>[r]^{\alpha_1}
  \ar[d]_{g_1}
  &
  B\ar@<2.00pt>[l]^{\beta_1}
  \ar@{=}[d]\\
  Z
  \ar[r]_{\sigma}
  &
  C
  \ar@<2.00pt>[r]^{\gamma}
  &
  B
  \ar@<2.00pt>[l]^{\delta}
  \\
  X
  \ar[r]_{\kappa_2}
  \ar[u]^{f}
  &
  A_2
  \ar[u]^{g_2}
  \ar@<2.00pt>[r]^{\alpha_2}
  &
  B\ar@{=}[u]
  \ar@<2.00pt>[l]^{\beta_2}
 }
}
 \end{equation}
if $\comm{1_X}{1_X}$ exists, then the unique (up to isomorphism) liftings of the split
extensions at the top and bottom of
\eqref{diagram:morphisms_of_split_exts_with_the_same_morphism_at_kernel}
along $\comm{1_X}{1_X}$ are isomorphic.
\end{proposition}
\begin{proof}
 We begin by forming the pullback
 \begin{equation}
  \vcenter{
   \label{diagram:pullback}
\xymatrix@!@C=-1ex@R=-2ex{
 X\times_Z X
  \ar[dd]_{\pi_1}
  \ar[rr]^-{{\kappa_1\times \kappa_2}}
  \ar[rd]^{\pi_2}&&
{A_1\times_C A_2}
 \ar[dd]^(0.30){{\pi_1}}|\hole
\ar@<2.00pt>[rr]^{{\alpha_1\pi_1}}
\ar[rd]^{{\pi_2}}&&
B
 \ar@{=}[dd]|(0.48)\hole|(0.51)\hole
 \ar@<2.00pt>[ll]^(0.4){{\langle \beta_1,\beta_2\rangle}}\ar@{=}[rd]&\\
&X\ar[rr]^(0.30){\kappa_2}\ar[dd]^(0.25){f}&&
A_2\ar[dd]^(0.3){g_2}\ar@<2.00pt>[rr]^(0.30){\alpha_2}&&
B\ar@{=}[dd]\ar@<2.00pt>[ll]^(0.70){\beta_2}\\
X\ar[rr]^(0.30){\kappa_1}|\hole\ar[rd]_{f}&&
A_1\ar@<2.00pt>[rr]^(0.30){\alpha_1}|\hole\ar[rd]_{g_2}&&
B\ar@<2.00pt>[ll]|\hole^(0.70){\beta_1}\ar@{=}[rd]&\\
&Z
\ar[rr]_{\sigma}&&
C
\ar@<2.00pt>[rr]^{\gamma}&&
B
\ar@<2.00pt>[ll]^{\delta}
}
 }
\end{equation}
in $\SplExt(\C)$.
Since by Proposition
 \ref{proposition:monomorphic behavior of central extensions},
 $\langle \comm{1_X}{1_X},\comm{1_X}{1_X}\rangle$ is the 
 commutator morphism of $1_{X\times_Z X}$ and itself, it follows
 from \wnh{} that
 $(\kappa_1\times\kappa_2) \langle \comm{1_X}{1_X},\comm{1_X}{1_X}\rangle$
 is normal. Therefore, Lemma \ref{lemma:lifting} produces a morphism
 \[
  \xymatrix{
 \como{X}{X}
 \ar[d]_{\langle\comm{1_X}{1_X},\comm{1_X}{1_X}\rangle}\ar[r]^-{{\kappa}}
 &
{A}\ar[d]^{{u_1}}
 \ar@<2.00pt>[r]^{{\alpha}}
 &
B
 \ar@{=}[d]
\ar@<2.00pt>[l]^{{\beta}}
 \\
 X\times_Z X
  \ar[r]_-{{\kappa_1\times \kappa_2}}
  &
{A_1\times_C A_2}
 \ar@<2.00pt>[r]^-{\alpha_1\pi_1}&
B,
 \ar@<2.00pt>[l]^-{\langle \beta_1,\beta_2\rangle}
}
 \]
 which after composing with the morphisms forming the upper and back faces
 of
 \eqref{diagram:pullback}
 produce liftings of the split extensions at the top
 and bottom of
 \eqref{diagram:morphisms_of_split_exts_with_the_same_morphism_at_kernel}
with same domain, as desired. 
\end{proof}
For a semi-abelian category $\C$ and an object $B$ in $\C$ let us
recall from \cite{BOURN_JANELIDZE:1998} that
the functor sending a split extension in $P^{-1}(B)$ (the fiber of
$P:\SplExt(\C)\to \C$ above $B$) to its kernel
turns out to be monadic, and the algebras of this monad, whose underlying
functor we denote by $B\flat -$, are called internal $B$-actions.
The induced equivalence associates to each split extension
\eqref{diagram:split}
the pair $(X,\chi)$, where $\chi$ is
the unique morphism making the diagram
\[
\xymatrix{
B\flat X
\ar[r]^{k}
\ar[d]_{\chi}
&
B+X
\ar@<0.5ex>[r]^-{[1,0]}
\ar[d]_{[\beta,\kappa]}
&
B
\ar@<0.5ex>[l]^-{\iota_1}
\ar@{=}[d]
\\
X
\ar[r]^{\kappa}
&
A\ar@<0.5ex>[r]^{\alpha}
&
B,
\ar@<0.5ex>[l]^{\beta}
\\
}
\]
in which $k$ is the kernel of $[1,0]$, commute. In the case where $\C$
is the category of groups,
it turns out that $B\flat X = \coprod_{b\in B} X$ and the internal $B$-actions
are in one to one correspondence with the usual $B$-actions.  Indeed, given $\chi$
as above for each $b$ in $B$ and each $x$ in $X$ defining $b\cdot x = \chi(\iota_b(x))$ - where
$\iota_b : X\to \coprod_{b\in B} X$ is a coproduct injection -
produces a $B$-action of $B$ on $X$. Translating Proposition
\ref{proposition_similtanius_lifting}
via this equivalence we obtain:
\begin{proposition}\label{proposition:actions}
Let $\C$ be a semi-abelian category satisfying \wnh. If $f : X\to Z$ is a morphism
in $\C$ with kernel central, and $\chi_1,\chi_2: B\flat X \to X$ and $\zeta: B\flat Z\to Z$
are internal $B$-actions such that $f$ is an internal $B$-action morphism from both
$(X,\chi_1)$ and $(X,\chi_2)$ to $(Z,\zeta)$, then 
$\chi_1$ and $\chi_2$ restrict to the same internal $B$-action on $[X,X]$.
\end{proposition}
Recall that a normal monomorphism $k:K\to X$ in a pointed category $\C$
is characteristic (in
agreement with the definition in \cite{CIGOLI_MONTOLI:2015}) if for
each normal monomorphism $n:X\to Y$ the composite $nk$ is a normal monomorphism.
Recalling also that in the category of groups the center of each object exists
and is characteristic we can recover the
group theoretical fact mentioned just above Proposition
\ref{proposition_similtanius_lifting} as a special case of the following
proposition. 
\begin{proposition} \label{proposition:actions_2}
Let $\C$ be a semi-abelian category satisfying \wnh{} in which the center of
each object exists and is characteristic. 
If $f : X\to Z$ is a morphism
in $\C$ with kernel central and $\chi_1,\chi_2: B\flat X \to X$ 
are internal $B$-actions such that $f\chi_1=f\chi_2$, then 
$\chi_1$ and $\chi_2$ restrict to the same internal  $B$-action on $[X,X]$.
\end{proposition}
\begin{proof}
Without loss of generality we may assume that $f$ is a normal
epimorphism (since otherwise factorizing $f$ as $me$ where
$m$ is a monomorphism and $e$ is a normal epimorphism produces a normal
epimorphism
$e$ with same properties as $f$). Since the kernel of $f$ is central
it follows that it factors through $z_X$, the center of $X$, and
hence, $q:X\to Y$, the cokernel of $z_X$, factors through $f$.
Since $z_X$ is characteristic, it follows by Proposition 2.7 of
\cite{CIGOLI_MONTOLI:2015} that there is a unique morphism $\zeta : B\flat Y\to Y$
such that $q$ is an internal $B$-action morphism from $(X,\chi_1)$ to $(Y,\zeta)$.
Since $f\chi_1=f\chi_2$, it follows that $q\chi_1=q\chi_2$ and so $q$ is also a morphism
from $(X,\chi_2)$ to $(Y,\zeta)$. We conclude via the Proposition
\ref{proposition:actions}. 
\end{proof}
\begin{remark}
It seems worth mentioning that if a category $\C$ has centralizers of
normal monomorphisms which are normal, then centers exist and are
characteristic. In such a category one easily proves that if $k:X\to Y$ is
normal monomorphism, then $k\wedge z_k \cong kz_X$.
\end{remark}
\section{Split extension classifiers of perfect objects}
In this section we state and prove our main results.

Recall that a reflexive graph $(A,B,\alpha_1,\alpha_2,\beta)$ in a category
$\C$ is a quadruple where $A$ and $B$ are objects in $\C$, and
$\alpha_1, \alpha_2: A\to B$ and $\beta : B\to A$ are morphisms in $\C$
such that $\alpha_1\beta = 1_B=\alpha_2\beta$. Recall also that a split
extension \eqref{diagram:split} is called faithful if there is at most
one morphism in $K^{-1}(X)$ from any split extension in $K^{-1}(X)$.
We have:
\begin{theorem}
\label{main_theorem}
 Let $\C$ be a pointed protomodular category satisfying \wnh{},
  let $(A,B,\alpha_1,\alpha_2,\beta)$ be a reflexive graph in $\C$, and let
$\kappa:X\to A$ be the kernel of $\alpha_1$. 
\begin{enumerate}[(a)]
\item \label{main_1} If a morphism $g:Y\to B$ commutes
with $\alpha_2 \kappa$ and
$\comm{1_X}{1_X}$ exists, then
 $\kappa \comm{1_X}{1_X}$ and $\beta g$ commute.
\item \label{main_2}If $X$ is perfect and the split extension 
\[
\xymatrix{
X \ar[r]^{\kappa} & A\ar@<0.5ex>[r]^{\alpha_1} & B\ar@<0.5ex>[l]^{\beta}
}
\]
is faithful, then the centralizer of $\alpha_2 \kappa$ is trivial.
\end{enumerate}
\end{theorem}

\begin{proof}
Suppose $\comm{1_X}{1_X}$ exists and that $g:Y\to B$ is a morphism commuting with $f=\alpha_2\kappa$. Consider the diagram
\begin{equation}
\label{big_diagram}
\vcenter{
\xymatrix@!@C=-4ex@R=-4ex{
 &&\como{X}{X}
 \ar[ld]_{\comm{1_X}{1_X}}
 \ar[ddd]^(0.18){\comm{1_X}{1_X}}|(0.33){\hole}|(0.67){\hole}
\ar[rrr]^-{\langle 0,1\rangle}
&&&
 Y\times \como{X}{X}
\ar@<0.5ex>[rrr]^{\pi_1}
\ar@{-->}[ld]_{u}
 \ar[ddd]^(0.18){1\times \comm{1_X}{1_X}}
|(0.3){\hole}
|(0.36){\hole}
|(0.64){\hole}
|(0.7){\hole}
&&&
Y
\ar@{=}[ld]
\ar@{=}[ddd]
\ar@<0.5ex>[lll]^{\langle 1,0\rangle}
\\
&X
\ar@{=}[ld]
\ar[rrr]^{\tilde \kappa}
\ar[ddd]^(0.18){f}|(0.33){\hole}
&&&
\tilde A
\ar@<0.5ex>[rrr]^(0.67){\tilde \alpha}
\ar[ddd]^(0.16){\langle\tilde \alpha,\alpha_2 h\rangle}|(0.3){\hole}|(0.36){\hole}
\ar[ld]_{h}
&&&
Y
\ar[ld]^(0.6){g}
\ar@{=}[ddd]
\ar@<0.5ex>[lll]^(0.33){\tilde \beta}
&
\\
X
\ar[ddd]_{f}
\ar[rrr]^{\kappa}
&&&
A
\ar@<0.5ex>[rrr]^{\alpha_1}
 \ar[ddd]_(0.45){\langle \alpha_1,\alpha_2\rangle}
&&&
B
\ar@{=}[ddd]
\ar@<0.5ex>[lll]^{\beta}
&&
\\
&&
X
\ar[ld]^{f}
\ar[rrr]|(0.33){\hole}^{\langle 0,1\rangle}|(0.67){\hole}
&&&
Y\times X
\ar@<0.5ex>[rrr]|(0.33){\hole}^{\pi_1}|(0.67){\hole}
\ar[ld]^(0.4){\langle \pi_1,\varphi\rangle}
&&&
Y
\ar@{=}[ld]
\ar@<0.5ex>[lll]|(0.33){\hole}^{\langle 1,0\rangle}|(0.67){\hole}
\\
&
B
\ar@{=}[ld]
\ar[rrr]^(0.45){\langle 0,1\rangle}|(0.67){\hole}
&&&
Y\times B
\ar@<0.5ex>[rrr]^{\pi_1}|(0.67){\hole}
\ar[ld]^(0.3){g\times 1}
&&&
Y
\ar[ld]^{g}
\ar@<0.5ex>[lll]|(0.33){\hole}^{\langle 1,g\rangle}
&
\\
B
\ar[rrr]^{\langle 0,1\rangle}
&&&
B\times B
\ar@<0.5ex>[rrr]^{\pi_1}
&&&
B
\ar@<0.5ex>[lll]^{\langle 1,1\rangle}
}
}
\end{equation}
in which
\begin{description}
 \item{-} $\varphi$ is the cooperator of $f$ and $g$;
\item{-} the object $(\tilde A,h,\tilde \alpha)$ is the pullback of
 $\alpha_1$ and $g$;
\item{-} the morphisms $\tilde \kappa$ and $\tilde \beta$ are the unique
 morphisms such that $h\tilde \kappa = \kappa$, $\alpha \tilde \kappa=0$,
  $h\tilde \beta = \beta g$
  and $\tilde \alpha \tilde \beta=1_Y$.
\end{description}
According to Proposition \ref{proposition_similtanius_lifting} (via
Lemma \ref{lemma:lifting}) there must be a
morphism $u:Y\times [X,X]\to \tilde A$ such that Diagram \eqref{big_diagram} consists of morphisms of split extensions.
The proof of \eqref{main_1} is completed by noting that $hu$ produces
a cooperator for $\kappa \comm{1_X}{1_X}$ and $\beta g$.
To prove \eqref{main_2}, we note that \eqref{main_1} tells us that
for any $g$ commuting with $f$, the morphisms $\kappa \comm{1_X}{1_X}$ and $\beta g$ commute.
Therefore, when $X$ is perfect $\kappa$ and $\beta g$ commute,
since $\comm{1_X}{1_X}$ is an isomorphism. But according to Proposition 5.2 of
\cite{BOURN_JANELIDZE:2009} (see also  Lemma 2.4 of \cite{CIGOLI_MANTOVANI:2012})
this means that $g=0$.
\end{proof}
\begin{remark}  
Recall that a split extension \eqref{diagram:split} can be equivalently defined
to be eccentric \cite{BOURN:2013b}, if every morphism $g:Y\to B$, such that
$\beta g$ and $\kappa$ commute, is a zero morphism. Noting that, the above
mentioned, Proposition 5.2 of \cite{BOURN_JANELIDZE:2009} which, using this language,
shows that ``faithful'' implies ``eccentric'', one can generalize (b) of
Theorem \ref{main_theorem} by replacing
``faithful'' with ``eccentric''.
\end{remark}
Suppose that $X$ is an object, in a pointed protomodular category,
such that $\spexc{X}$ exists.
Since the generic split extension \eqref{diagram:generic_split} is
necessarily faithful and (as recalled above) forms part of a
groupoid with underlying reflexive graph $(\spexc{X}\ltimes X, p_1,p_2, i)$
such that $c_X=p_2k$,
the previous theorem produces the following corollaries.
\begin{corollary}\label{corollary:a}
Let $\C$ be a pointed protomodular category in satisfying \wnh{}.  If $X$ is an object such that $\comm{1_X}{1_x}$,
$\spexc{X}$ and $z_{c_X}$
exists, then the morphisms $k \comm{1_X}{1_X}$ and $i z_{c_X}$ commute.
\end{corollary}
\begin{corollary}\label{corollary:b}
Let $\C$ be a pointed protomodular category satisfying \wnh{}.
If $X$ is a perfect object in $\C$ such that $\spexc{X}$ exists,
 then $z_{c_X}=0$ and hence $z_{\spexc{X}}=0$.
\end{corollary}

\begin{remark}
Corollary \ref{corollary:a} applied to the category of groups
shows that if  $\theta : X\to X$ is a central automorphism, then
$\theta([x,y]) = [x,y]$.
Applied to the category of Lie algebras it shows that if $d$ is a central
derivation, then $d([x,y])=0$.
\end{remark}
As an immediate consequence of Corollary \ref{corollary:b} we obtain:
\begin{corollary}
Let $\C$ be a pointed protomodular category admitting all generic split
 extensions and satisfying \wnh{}.
If $X$ is a perfect object in $\C$, then $z_{c_X}=0$ and hence $z_{\spexc{X}}=0$.
\end{corollary}
\begin{example}
Recall from \cite{CIGOLI_GRAY_VAN_DER_LINDEN:2015b} that \emph{algebraic
coherence} implies \nh{} which is stronger than \wnh{}. Recall also that there
are plenty of algebraically coherent categories, including
all categories of interest in the sense of G. Orzech \cite{ORZECH:1972}.
It follows that the above results apply to all action representable
algebraically coherent categories. These include
the categories of groups, Lie algebras, boolean rings and cocommutative Hopf
algebras (for action representability of the first three categories see
\cite{BORCEUX_JANELIDZE_KELLY:2005b} and for the last one see 
\cite{GRAN_STERCK_VERCRUYSSE:2019},
for algebraic coherence see 
\cite{CIGOLI_GRAY_VAN_DER_LINDEN:2015b}).
Further examples can be obtained by taking a functor category of, or
the category of internal groupoids in,
a semi-abelian, algebraically coherent category with split extension
classifiers and normalizers (see \cite{GRAY:2013b} and \cite{GRAN_GRAY:2021}).
This produces, amongst many others,
crossed modules of groups and Lie algebras as further examples.
\end{example}
\begin{proposition}
If $\mathcal{V}$ is a variety of semi-abelian algebras satisfying \wnh{}, then 
the category $\mathcal{V}(\mathbf{Top})$ of internal $\mathcal{V}$ algebras in
the category of topological spaces satisfies \wnh{}.
\end{proposition}
\begin{proof}
Recall that coequalizers and equalizers are calculated as in $\mathcal{V}$, equipped with
the quotient and subspace topology, respectively (see 
\cite{BORCEUX_CLEMENTINO:2005}). This means that
$\comm{1_X}{1_X}= \ke(\coke(\langle 1_X,1_X\rangle) \langle 1,0\rangle)$
is calculated as in $\mathcal{V}$ with the subspace topology. Since a subspace 
of a subspace is a subspace of the whole sapce, this also implies
that a normal monomorphism in $\mathcal{V}(\mathbf{Top})$ is characteristic
if and only if it is characteristic in $\mathcal{V}$.
\end{proof}
\begin{example}
The category of topological groups has been shown to be action
representable \cite{CAGLIARI_CLEMENTINO:2019} and satisfies \wnh{} by the previous proposition.
It follows that our results hold for the category of topological groups.
\end{example}
\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}